marks test scores were 87, 87, 67,88,78 . what does he need on his next test to average 80?
Answer by Alan3354(69443)

marks test scores were 87, 87, 67,88,78 . what does he need on his next test to average 80?
------
An 80 average for 6 tests --> 6*80 = 480 points.
---
87 + 87 + 67 + 88 + 78 = 407
480 - 407 = 73 needed
=======================
What's the highest average he can make in the class?